OpenStudy (anonymous):

what is the domain and the range of f(x)=1/x+1

OpenStudy (anonymous):

get zero's of this function to get domain. then get its inverse and get its domain thus you will get is range

OpenStudy (anonymous):

?

OpenStudy (anonymous):

why are you sticking with that? domain is all numbers except which make the function undefined.

OpenStudy (anonymous):

is it all real numbers except for 1?

OpenStudy (kainui):

Are you asking if the domain or range is all real numbers except for 1? And why do you think that?

OpenStudy (anonymous):

is the domain all real numbers except for 1?

OpenStudy (freckles):

your function is 1/(x+1) right? if x=1, then you have 1/(1+1)=1/2 is 1/2 not a number?

OpenStudy (freckles):

I think you are thinking about the opposite of 1 -1 (x+1) is what you don't want to be zero since you can not divide by zero and we know x+1=0 when x=-1

OpenStudy (freckles):

the range is kinda easy to find here too we have 1/(x+1) hint: we know the only thing that will ever make a fraction 0 is the top